# Federated Continual Learning Goes Online: Uncertainty-Aware Memory Management for Vision Tasks and Beyond

### v3

@@ 5 Discussion and Conclusion
Practicality and computational time - Appendices A.6 and A.7. In terms of practicality, for standard datasets like CIFAR10 and CIFAR100, the number of communication rounds per task with a jump parameter is comparable to the one reported in FedCIL (40 communication rounds per task) and considerably smaller than MFCL (100 communication rounds per task). The same consideration is also valid for the training computational time; our approach has a comparable computational time compared to FedCIL and a considerably smaller one compared to MFCL.
@@
Limitations. A limitation of the proposed approach is the need to use an ensembling technique (in our case, test-time augmentation) in order to compute the BI-based estimates. This may result in a reduced efficiency compared to, e.g., ER. Furthermore, the TTA-based estimation of the Bregman Information is only an estimate of the true unknown uncertainty; the estimator used throughout the experiments (Eq. 1) is only asymptotically unbiased and underestimates the theoretical quantity (Gruber and Buettner, 2023).
@@ A.1 Predictive Uncertainty Scores
Least Confidence (LC) [Culotta and McCallum, 2005] measures the predictive uncertainty by looking at the samples with the smallest predicted class probability. If the probability associated with the most probable class is low, the model is less certain about the given sample.
@@
Margin Sampling (MS) [Campbell et al., 2000] measures the predictive uncertainty by looking at the difference between the most probable predicted class and the second largest one . If the two probabilities are similar, the model is uncertain.
@@
Ratio of Confidence (RC) [Campbell et al., 2000] is similar to MS. In this case, instead of computing the difference, the ratio between the probabilities of the two most probable classes is considered.
@@
Entropy (EN) [Shannon, 1948] considers, differently from the previously introduced metrics, the whole probability distribution. The entropy is computed as follows:
@@ A.2 Set of Augmentations
{minted}python transform_cands = [ CutoutAfterToTensor(args, 1, 10), CutoutAfterToTensor(args, 1, 20), v2.RandomHorizontalFlip(), v2.RandomVerticalFlip(), v2.RandomRotation(degrees=10), v2.RandomRotation(45), v2.RandomRotation(90), v2.ColorJitter(brightness=0.1), v2.RandomPerspective(), v2.RandomAffine(degrees=20, translate=(0.1, 0.3), scale=(0.5, 0.75)), v2.RandomResizedCrop(args.input_size[1:], scale=(0.8, 1.0), ratio=(0.9, 1.1), antialias=True), v2.RandomInvert() ]
@@ A.3 Dataset Statistics
In Table 6, we report the statistics of the datasets used for the main experiments. We use dataset from different domains (CIFAR10, CIFAR100, CRC-Tissue, KC-Cell) and with different modalities (20NewsGroups, Yahoo Answers, DBPedia). For CRC-Tissue, in order to create tasks with decreasing size and equal number of classes, we remove the smallest class from our evaluation. For DBPedia and Yahoo Answers Zhang et al. [2015], we randomly select a subset of 25000 and 33600 samples respectively.
@@
A.4 Evaluation Metrics
In line with Yoon et al. [2021a] and Wuerkaixi et al. [2024], we employ the average Last Accuracy (A) and average Last Forgetting (F) – a federated adaptation of the metrics defined in Chaudhry et al. [2018a]. Last refers to the measurement of the value at the end of the stream for all the clients. Suppose represents the accuracy of task after learning task on client . The last accuracy at task on client is defined as . Let represent the total number of clients. The average last accuracy is then defined as:
@@ A.4 Evaluation Metrics
A.5 Ablation Study
In this section, we investigate the effect of various hyperparameters (i.e., burn-in period, jump parameter, parameter averaging strategy, and set of augmentations) on the learning capabilities of the proposed approach while also providing insights and explanations for best practices to use them in this newly proposed context.
@@ A.5 Ablation Study
Parameter averaging strategy. The parameter averaging strategy can be changed without affecting particularly the performance of the approach as can be seen in Table 8. Overall, taking the average between the current and the previous global parameters (as proposed in Shenaj et al. [2023]) helps in reducing CF compared to the vanilla FedAvg.
@@
A.6 Practicality Analysis - Number of Communication Rounds
@@
A.7 Time Complexity Analysis
@@
A.8 Performance Analysis of Generative-based Models
The proposed online-FCL (O-FCL) scenario is very different from the standard FCL scenario presented in, e.g., MFCL [Babakniya et al., 2024] and related works based on generative-based solutions. In the standard scenario, they assume to work on task datasets that can be iterated multiple times. For instance, let us consider MFCL on CIFAR100 in which the authors assume to have 100 communication rounds for each task. This implies that the whole task dataset is processed 100 times. Thus, after many iterations over the same task dataset, the generative model can effectively generate synthetic images for the given task. In our online scenario, instead, we only have access to one batch of 10 images at every iteration. In this setting, generative-based solutions are limited in their capabilities as they require many data points and many iterations over the training data for being effective. For this reason, we believe that the performance gap compared to the standard FCL setting is given by the inefficiency of generative models in online settings where we do not have the possibility to retrain the model on the same data multiple times. This is in line with the literature in online-CL where the most successful solutions rely on memory-based solutions [Ma et al., 2022, Soutif-Cormerais et al., 2023]. To prove this point, we generated some images using MFCL at the end of the training procedure in the online setting (see Figure 5). Differently from the images reported in the original paper, we can see that no informative pattern has been discovered; the synthetic images provide limited information about the past information. This is also corroborated by the results of MFCL on the datasets for vision tasks reported in the main paper. From the results, we can see that MFCL is better than standard approaches without memory, but it is not competitive compared to memory-based approaches.

### v4

@@ 5 Discussion and Conclusion
Practicality and computational time - Appendices A.8 and A.9. In terms of practicality, for standard datasets like CIFAR10 and CIFAR100, the number of communication rounds per task with a jump parameter is comparable to the one reported in FedCIL (40 communication rounds per task) and considerably smaller than MFCL (100 communication rounds per task). The same consideration is also valid for the training computational time; our approach has a comparable computational time compared to FedCIL and a considerably smaller one compared to MFCL.
@@
Limitations. A limitation of the proposed approach is the need to use an ensembling technique (in our case, test-time augmentation) in order to compute the BI-based estimates. This may result in a reduced efficiency compared to, e.g., ER. Furthermore, the TTA-based estimation of the Bregman Information is only an estimate of the true unknown uncertainty; the estimator used throughout the experiments (Eq. 1) is only asymptotically unbiased and underestimates the theoretical quantity (Gruber & Buettner, 2023).
@@ Acknowledgments
Reproducibility Statement
Code availability. The code used to implement our approach and to conduct the experiments is available at https://github.com/MLO-lab/online-FCL.
Data accessibility. All the datasets employed in the experiments are publicly available. We provide detailed instructions on how we use them. In case some particular preprocessing step is required, the corresponding scripts are included in our code repository.
Hyperparameters. All hyperparameters and training details are described in the main paper. The effect of key hyperparameter choices is investigated in the ablation study reported in Appendix A.7.
Hardware and runtime. As mentioned in the main paper, our experiments were conducted on a Linux machine using a single Quadro RTX 5000 and 16 GB RAM. In Tables 14 and 15, we report the expected training runtime for our approach and the mini-batch processing time in seconds respectively.
Experiment instructions. The code repository includes a README file with instructions to run the experiments. Furthermore, to facilitate the comprehension of our approach, Appendix A.11 describes the pseudocode of the proposed methodology.
@@ A.1 Predictive Uncertainty Scores
Least Confidence (LC) (Culotta & McCallum, 2005) measures the predictive uncertainty by looking at the samples with the smallest predicted class probability. If the probability associated with the most probable class is low, the model is less certain about the given sample.
@@
Margin Sampling (MS) (Campbell et al., 2000) measures the predictive uncertainty by looking at the difference between the most probable predicted class and the second largest one . If the two probabilities are similar, the model is uncertain.
@@
Ratio of Confidence (RC) (Campbell et al., 2000) is similar to MS. In this case, instead of computing the difference, the ratio between the probabilities of the two most probable classes is considered.
@@
Entropy (EN) (Shannon, 1948) considers, differently from the previously introduced metrics, the whole probability distribution. The entropy is computed as follows:
@@ A.3 Dataset Statistics
In Table 6, we report the statistics of the datasets used for the main experiments. We use dataset from different domains (CIFAR10, CIFAR100, CRC-Tissue, KC-Cell) and with different modalities (20NewsGroups, Yahoo Answers, DBPedia). For CRC-Tissue, in order to create tasks with decreasing size and equal number of classes, we remove the smallest class from our evaluation. For DBPedia and Yahoo Answers Zhang et al. (2015), we randomly select a subset of 25000 and 33600 samples respectively.
@@
A.4 Additional results on TinyImageNet
Compared to CIFAR datasets, TinyImageNet (Le & Yang, 2015) consists of more classes (200) with larger images ( instead of ), providing a slightly different scenario compared to previously explored datasets. In Table 7, we report the results for all the baselines considered in our experiments (except FedCIL which does not consider images larger than ). Similar to the empirical findings reported in the main paper, BI is able to maintain competitive predictive performance while having the smallest forgetting value.
| A() | F() | |
| ER | 11.26 1.31 | 6.31 1.23 |
| CBR | 11.06 1.65 | 6.44 2.08 |
| LC | 10.16 0.63 | 6.96 1.20 |
| MS | 10.84 1.08 | 6.27 1.13 |
| RC | 10.88 1.08 | 6.76 1.17 |
| EN | 10.18 1.02 | 6.69 2.84 |
| FedAvg | 1.74 0.30 | 12.49 1.14 |
| W. FedAvg | 1.20 0.29 | 10.96 0.74 |
| FedProx | 1.50 0.25 | 10.98 0.51 |
| FedCIL | ✗ | ✗ |
| MFCL | 3.60 0.49 | 21.58 0.60 |
| BI (ours) | 11.04 0.35 | 6.07 0.74 |
A.5 Evaluation Metrics
In line with Yoon et al. (2021a) and Wuerkaixi et al. (2024), we employ the average Last Accuracy (A) and average Last Forgetting (F) – a federated adaptation of the metrics defined in Chaudhry et al. (2018a). Last refers to the measurement of the value at the end of the stream for all the clients. Suppose represents the accuracy of task after learning task on client . The last accuracy at task on client is defined as . Let represent the total number of clients. The average last accuracy is then defined as:
@@ A.5 Evaluation Metrics
A.6 Clarification about Figure 3
The notable performance gap on the last task graphically describes what is the effect of catastrophic forgetting (CF) on the learning model. One can see how approaches without replay mechanisms (such as FedAvg, the dashed blue line in the figure) mainly focus on the last task with very good performance ( accuracy) while completely forgetting the knowledge of the previous ones ( accuracy). This effect can be mitigated with generative-based replay mechanisms (e.g., MFCL). In our case, thanks to the proposed approach, the accuracy across tasks is kept more uniform (between and ), leading to better average accuracy and reduced CF.
A.7 Ablation Study
In this section, we investigate the effect of various hyperparameters (i.e., burn-in period, jump parameter, parameter averaging strategy, set of augmentations, and number of tasks) on the learning capabilities of the proposed approach while also providing insights and explanations for best practices to use them in this newly proposed context.
@@ A.7 Ablation Study
Parameter averaging strategy. The parameter averaging strategy can be changed without affecting particularly the performance of the approach as can be seen in Table 9. Overall, taking the average between the current and the previous global parameters (as proposed in Shenaj et al. (2023)) helps in reducing CF compared to the vanilla FedAvg.
@@
Number of tasks. For this experiment, we consider CIFAR100 as it allows us to create a larger number of tasks. From the results in Table 12, the increasing number of tasks seems to mainly influence catastrophic forgetting in all considered cases. The results look reasonable since, by increasing the number of tasks, it is more difficult to retain knowledge of the initial tasks. The same trend is also observed for the baselines included for comparison, with BI consistently outperforming ER and MFCL, with large gains especially over ER for large task numbers.
| Tasks | 5 | 10 | 20 | 25 | |
| ER | A | 10.95 | 8.59 | 9.43 | 9.09 |
| F | 9.12 | 20.06 | 26.76 | 41.56 | |
| MFCL | A | 4.60 | 4.43 | 3.98 | 3.58 |
| F | 13.69 | 18.59 | 31.60 | 37.43 | |
| BI (ours) | A | 14.31 | 14.56 | 14.46 | 13.71 |
| F | 8.96 | 14.36 | 19.94 | 33.65 |
A.8 Practicality Analysis - Number of Communication Rounds
@@
A.9 Time Complexity Analysis
@@
In the online scenario, the total number of samples in a dataset does not affect the scalability of an algorithm. In fact, since the method does not use the whole dataset at once but rather process mini-batches sequentially, the only important factor is the time for processing a mini-batch. For this, an important factor that may affect the computational time is the image size. Table 15 reports the effect of the image size on the time for processing a mini-batch.
| Image size | BI (ours) | MFCL |
| 32 32 | 0.20 | 0.40 |
| 64 64 | 0.42 | 0.75 |
| 128 128 | 1.32 | ✗ |
| 224 224 | 4.19 | ✗ |
A.10 Performance Analysis of Generative-based Models
The proposed online-FCL (O-FCL) scenario is very different from the standard FCL scenario presented in, e.g., MFCL (Babakniya et al., 2024) and related works based on generative-based solutions. In the standard scenario, they assume to work on task datasets that can be iterated multiple times. For instance, let us consider MFCL on CIFAR100 in which the authors assume to have 100 communication rounds for each task. This implies that the whole task dataset is processed 100 times. Thus, after many iterations over the same task dataset, the generative model can effectively generate synthetic images for the given task. In our online scenario, instead, we only have access to one batch of 10 images at every iteration. In this setting, generative-based solutions are limited in their capabilities as they require many data points and many iterations over the training data for being effective. For this reason, we believe that the performance gap compared to the standard FCL setting is given by the inefficiency of generative models in online settings where we do not have the possibility to retrain the model on the same data multiple times. This is in line with the literature in online-CL where the most successful solutions rely on memory-based solutions (Ma et al., 2022; Soutif-Cormerais et al., 2023). To prove this point, we generated some images using MFCL at the end of the training procedure in the online setting (see Figure 5). Differently from the images reported in the original paper, we can see that no informative pattern has been discovered; the synthetic images provide limited information about the past information. This is also corroborated by the results of MFCL on the datasets for vision tasks reported in the main paper. From the results, we can see that MFCL is better than standard approaches without memory, but it is not competitive compared to memory-based approaches.